What is the probability that at most 40 boxes will be delivered on a particular day?


1
Expert's answer
2021-05-24T03:47:15-0400

The probability distribution for the question is as shown below:



The probability that at most 40 boxes will be delivered on a particular day will be calculated as:

"P\left(X\le 40\right)"

So the value of X are "P(40), P(37) and P(35)."

"P\left(X\le \:40\right)=P\left(40\right)+P\left(37\right)+P\left(35\right)\:"

"P\left(X\le \:\:\:40\right)=\frac{1}{10}+\frac{1}{5}+\frac{1}{10}=\frac{4}{10}\:=\frac{2}{5}\:or\:0.40"